OpenStudy (anonymous):

Describe the differences between the graph of y-3/4(x-6)^2+4 and the standard position graph of y = x^2.

OpenStudy (anonymous):

y = x^2 is the "parent function"... right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

graph of y = 3/4(x-6)^2 + 4 ?

OpenStudy (anonymous):

(In your question, there is no equal sign in the first equation)

OpenStudy (anonymous):

it's y = -3/4(x-6)^2 + 4

OpenStudy (anonymous):

srry

OpenStudy (anonymous):

OK... There are four components to graphing a quadratic...

OpenStudy (anonymous):

First, the "vertical shift" .... In this case, + 4....

OpenStudy (anonymous):

Second, the "horizontal shift" --- movement along the x-axis... In this case, (x - 6) implies a horizontal shift to the RIGHT...

OpenStudy (anonymous):

And, third... the "stretch" is 3/4....

OpenStudy (anonymous):

isn't it -3/4?

OpenStudy (anonymous):